OpenStudy (anonymous):

Which ordered pair is a solution for this system of inequalities? 2x + 3y > 1 and y < 5/4 x - 1 Answer (3, 9) (-4, 2) (8, -1) (-2, -5)

OpenStudy (anonymous):

so 2x+3y > 1, y < 5x-4/4 2x+3y>1, 4y<5x-4 Now you have to use the values in the options. On using, (8,-1) 2x8 + 3x-1 = 16-3 = 13, which is > 1. 4y < 5x-4 4.-1 < 5.8 - 4 => -4 < 36, which is true. so (8,-1) is the answer.